It is estimated

30

% of all adults in United States invest in stocks and that

84

% of U.S. adults have investments in fixed income instruments (savings accounts, bonds, etc.). It is also estimated that

26

% of U.S. adults have investments in both stocks and fixed income instruments.

(a) What is the probability that a randomly chosen U.S. adult invests in stocks, given that s/he invests in fixed income instruments? Round your answer to

2

decimal places.

(b) What is the probability that a randomly chosen stock investor also invests in fixed income instruments? Round your answer to

2

decimal places.
Answer #1

p(stock)=0.30

p(income instrument)=0.84

p(both)=0.26

a)P(Stocks/income)=(0.27)/(0.84)=0.3214

=0.32

b)p(income/stocks)=p(both)/p(stocks)

=0.27/0.30

=0.9